Let u =< 3, −1, −2 >, v =< −1, 0, 2 > and w =< −6, 1, 4 >.

Compute the expressions below.

a) >v + 3w

b) 3>u − 2>v (1)

c) −(>v + 3>u). 


1
Expert's answer
2021-06-30T12:27:30-0400

a) "v+3w=<-1+(-6)*3, 0+1*3, 2+4*3>=<-19,3,14>."


b) "3u-2v=<3*3-(-1)*2,(-1)*3-0*2,(-2)*3-2*2>=<11,-3,-10>."


c) "-(v+3u)=-<-1+3*3,0+(-1)*3,2+(-2)*3>=<-8, 3, 4>."